# thanks: These two authors contributed equallythanks: These two authors contributed equally

### v3

@@ A.1.4 Gamma function and distribution conventions
There are a number of conflicting conventions for the gamma function and distribution arguments. We will use those defined on Wikipedia [37]. The gamma distributed random variable $X$ will be written:
@@ A.2 Methods: The derivation of the RLTO growth rate
A.2.1 The metabolic load of protein and the Hwa model
To model the effect of the metabolic load on cell growth, we will expand on a model used by Hwa and co-workers [14]. For conciseness, we will call the original model the Hwa model.
@@ A.2.1 The metabolic load of protein and the Hwa model
The key assumption in the Hwa model is that the abundances of the R and P sectors can change in size to accommodate changes in the nutrient quality and translation load associated with a particular growth condition [14]. In contrast, the Q sector has a fixed proteome fraction, $\Phi_{Q}$, irrespective of growth conditions. In the Hwa model, the size of these adjustable R and P sectors are chosen to optimize the growth rate $k$.
@@
Note that our generalized Hwa model is written for arbitrary number of functional sectors $i\neq Q$ and the key determinant of the change in growth rate is the fraction of productive protein $\Phi_{*}$. This equation will be used to model the fitness cost of the metabolic load.
@@ A.2.4 Growth rate with stochastic arrest
is the exponential mean, where $T$ is the stochastic cell cycle duration and $\mathbb{E}$ is the expectation operator [40, 29].
@@ C.2.2 Estimating the transcription rate: $\beta_{m}$
To estimate the transcription rate for gene $i$, we start from the estimated cellular message number $N_{m/c,i}$ and use the Paulsson model prediction for the cellular message number:
@@ C.2.3 Estimating the message number: $\mu_{m}$
To estimate the message number of gene $i$, we use the predicted value from the steady-state stochastic kinetic model (see A.1.1):
@@ Appendix E Data Tables
Data S5: Organism: S. cerevisiae (yeast). Original sources: Protein abundance [21]. mRNA abundance: [70]. Processing: We merged these datasets. We added protein number and message number (messages per cell-cycle) as described in Sec. C.
Data S6: Organism: Mus musculus (mammalian). Original source: [23]. Processing: We added protein number and message number (messages per cell-cycle) as described in Sec. C.
Data S7: Organism: E. coli. Original source: [22]. Processing: We added protein number and message number (messages per cell-cycle) as described in Sec. C.
@@
Data S8: Organism: S. cerevisiae (yeast). Original sources: Noise measurements: [8]. Essential gene classifications: [19]. Protein fluorescence data, rescaled by mass-spec data as described in Sec. D.4.1: [8]. Processing: We merged these datasets. We added protein number as described in Sec. C.

### v4

@@ A.1.4 Gamma function and distribution conventions
There are a number of conflicting conventions for the gamma function and distribution arguments. We will use those defined on Wikipedia and the CRC Encyclopedia of Mathematics [37]. The gamma distributed random variable $X$ will be written:
@@ A.2 Methods: The derivation of the RLTO growth rate
A.2.1 The metabolic load of protein and the resource allocation model
To model the effect of the metabolic load on cell growth, we will expand on a model used by Hwa and co-workers [14]. For conciseness, we will call the original model the resource allocation model.
@@ A.2.1 The metabolic load of protein and the resource allocation model
The key assumption in the resource allocation model is that the abundances of the R and P sectors can change in size to accommodate changes in the nutrient quality and translation load associated with a particular growth condition [14]. In contrast, the Q sector has a fixed proteome fraction, $\Phi_{Q}$, irrespective of growth conditions. In the resource allocation model, the size of these adjustable R and P sectors are chosen to optimize the growth rate $k$.
@@
Note that our generalized resource allocation model is written for arbitrary number of functional sectors $i\neq Q$ and the key determinant of the change in growth rate is the fraction of productive protein $\Phi_{*}$. This equation will be used to model the fitness cost of the metabolic load.
@@ A.2.4 Growth rate with stochastic arrest
is the exponential mean, where $T$ is the stochastic cell cycle duration and $\mathbb{E}$ is the expectation operator [40, 29]. We take a coarse-grained model which considers changes in growth rate due to fluctuations in protein number to be negligible. Note that Eq. 44 is equivalent to the Euler-Lotka equation [41, 42].
@@ C.2.2 Estimating the transcription rate: $\beta_{m}$
To estimate the transcription rate for gene $i$, we start from the estimated cellular message number $N_{m/c,i}$ and use the canonical steady-state noise model prediction for the cellular message number:
@@ C.2.3 Estimating the message number: $\mu_{m}$
To estimate the message number of gene $i$, we use the predicted value from the canonical steady-state noise model (see A.1.1):
@@ Appendix E Data Tables
Data S5: Load-balancing data for yeast cells. Organism: S. cerevisiae (yeast). Original sources: Protein abundance [21]. mRNA abundance: [72]. Processing: We merged these datasets. We added protein number and message number (messages per cell-cycle) as described in Sec. C.
Data S6: Load-balancing data for mammalian cells. Organism: Mus musculus (mammalian). Original source: [23]. Processing: We added protein number and message number (messages per cell-cycle) as described in Sec. C.
Data S7: Load-balancing data for E. coli. Organism: E. coli. Original source: [22]. Processing: We added protein number and message number (messages per cell-cycle) as described in Sec. C.
@@
Data S8: Noise data for yeast. Organism: S. cerevisiae (yeast). Original sources: Noise measurements: [8]. Essential gene classifications: [19]. Protein fluorescence data, rescaled by mass-spec data as described in Sec. D.4.1: [8]. Processing: We merged these datasets. We added protein number as described in Sec. C.
E.4 Datasets for one-message-rule exceptions.
Data S9: Exceptions to the one-message-rule for yeast cells. Organism: S. cerevisiae (yeast). Original sources: Message abundances: [72]. Essential gene classifications: [19].
Data S10: Exceptions to the one-message-rule for human cells. Organism: H. sapiens (human). Original sources: Message abundances: [62]. Essential gene classifications: [53].
Data S11: Exceptions to the one-message-rule for E. coligrown in rich media (LB). Organism: E. coli. Original sources: Message abundances: [46]. Essential gene classifications: [55].